# Observation of an Accreting Planetary-Mass Companion with Signs of Disk–Disk Interaction in Orion ††thanks: Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme ID 114.28JY

### v1

@@
Observation of an Accreting Planetary-Mass Companion with Signs of Disk–Disk Interaction in Orion ††thanks: Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme ID 114.28JY
Young ( Myr) planetary-mass companions (PMCs) provide valuable insights into the formation and early evolution of planetary systems. To date, only a dozen such objects have been identified through direct imaging. Using JWST/NIRCam observations towards the Orion Nebula, obtained as part of the PDRs4All Early Release Science program, we have identified a faint point source near the M-type star V2376 Ori. Follow-up spectroscopic observations with the MUSE instrument on the VLT confirm that the source, V2376 Ori b, is indeed a young planetary-mass companion. It is a member of Orion D, around 80 pc in the foreground of the Trapezium cluster of Orion and with an age of approximately Myr. We fit the SED of V2376 Ori b to infer a mass of . The MUSE spectrum reveals several accretion tracers. Based on the H line intensity, we estimate an accretion rate of 10, which is comparable to that of young PMCs such as PDS 70b. In addition, the MUSE data cube reveals extended emission in the [O ii] doublet at 7320 and 7330 Å, which is interpreted as evidence of a dynamical interaction between the two sources that, potentially, involves mass transfer between their individual accretion disks. These results demonstrate that JWST/NIRCam imaging surveys of young stellar associations can uncover new PMCs, which can then be confirmed and characterized through ground-based spectroscopic follow-up.
@@
brown dwarfs – exoplanets – stars: low-mass – accretion, accretion disks – direct imaging1 Introduction
Direct imaging has revealed key properties of planetary-mass companions (PMCs, , au; Currie et al. 2022). Only a dozen have been identified in young ( Myr) stellar clusters (e.g., Currie et al., 2013; Luhman et al., 2006; Schmidt et al., 2008; Bowler et al., 2014). Despite their rarity, young PMCs offer valuable constraints on planetary formation pathways (Kratter & Lodato, 2016; Schlaufman, 2018), especially in environments where external factors like photoevaporation and dynamical interactions may play a role (Reiter & Parker, 2022; Planet formation environments collaboration, 2025). Expanding the sample of known young PMCs is critical but has been limited by the time-intensive nature of ground-based high-contrast imaging. JWST has recently demonstrated its potential to image both known and new PMCs as well as exoplanets (Boccaletti et al., 2024; Matthews et al., 2024; Carter et al., 2023; Mullally et al., 2024; Lagrange et al., 2025). In this letter, we report the discovery of a 7 Myr-old PMC in Orion, identified in JWST/NIRCam data from the PDRs4All Early Release Science program (PID 1288; Berné et al., 2022; Habart et al., 2024), and confirmed via follow-up spectroscopy with VLT/MUSE.
@@ 2 Source and data
V2376 Ori (, ) is a young star situated northeast of the Orion Nebula Cluster (ONC), cataloged as an M-type star by Hillenbrand (1997). In their catalog of protoplanetary disks, Ricci et al. (2008) identified a faint source within of V2376 Ori, which the authors interpreted as a jet. This faint source was also reported by Habart et al. (2024) in the Paschen image obtained with the JWST. In these earlier studies, V2376 Ori was believed to be part of the ONC association. However, Gaia observations provide a parallax for V2376 Ori of mas, placing the source at a distance of pc, which is much closer than the ONC (390 pc, Maíz Apellániz et al., 2022). This implies that V2376 Ori is instead part of the Orion D association identified by Kounkel et al. (2018), a population of young stars (7-10 Myr) situated in front of the ONC. This foreground population has also been identified by Alves & Bouy (2012) as a region dominated by the NGC 1980 cluster, aged 4-5 Myr. Thus, we assume an age of Myr for V2376 Ori.
@@ 3 Results
Fig. 1 presents the HST and JWST images of V2376 Ori obtained in 12 filters. The faint source previously reported by Ricci et al. (2008) and Habart et al. (2024) is clearly detected in all bands. Fig. 2 shows the MUSE integral field spectroscopic observations of the system, it combines integrated maps and the spectrum of the companion over the full MUSE wavelength range. Panel 2A displays the white light image integrated over the 6908–9158 Å range. The companion source is clearly visible. Panel 2B shows the continuum-subtracted map (using the Background2D procedure described in Appendix C) of the forbidden [C i] 8727 line from neutral atomic carbon, where the companion is also detected. In addition, we find extended emission in the [O ii] doublet at 7320 and 7330 Å, which is shown in panel 2C as the sum of the integrated maps of both lines. A similar, although fainter, structure is also visible in the [O i] 6300 map. We analyze the relative astrometry of V2376 Ori and the nearby point source in Appendix B.1. We measure a angular separation of , corresponding to a projected separation of au when assuming a distance of pc. The relative position of the two objects has remained constant within instrumental uncertainties over 20 years, consistent with the source being gravitationally bound. We therefore identify it as a companion, hereafter referred to as V2376 Ori b. Panel 2D presents the MUSE spectrum of this companion.
3.1 Temperature, luminosity and mass of V2376 Ori b
Fig. 3 presents the HST and JWST spectral energy distribution (SED) of V2376 Ori b, obtained as explained in Appendix C, within the red dashed circular aperture of 0.08” radius, shown in Fig. 1 (G). We fit a black-body using a grid of temperatures ranging from 1000 K to 3000 K with a step of 100 K, and a scaling coefficient, with values ranging from to . The best-fit parameters are those that minimize the Root Mean Square Error (RMSE). The error bars are determined considering the parameters range for which the RMSE is within 1.5 times the minimum RMSE. This yields K. The derived black-body luminosity is , where the uncertainty arises from the errors on both the temperature and the distance to the source. As the system is situated in front of the ONC, we consider that extinction is negligible. Fig. 5 shows the position of V2376 Ori b on the evolutionary tracks from Baraffe et al. (2003). Considering the obtained luminosity, effective temperature, and an age of 73 Myr, this yields a mass in the range of 10-30 . The derived physical properties of V2376 Ori b are summarized in Table 1. Fig. 7 shows that these properties are similar to those found for other directly imaged PMCs and giant exoplanets.
@@ 3.2 MUSE spectroscopy of the V2376 Ori system
Fig. 8 present the MUSE spectrum of V2376 Ori (see details on the extraction method in Appendix E). This spectrum is found to be compatible with that of an M7.5-9 spectral type (Appendix F). Fig. 2 (D) presents the MUSE spectrum for V2376 Ori b. It shows smooth continuum emission, with an increase towards the red and numerous emission lines. This spectrum is similar to that of the MUSE spectrum of the 12 MJup planetary-mass companion Delorme 1 (AB) (Eriksson et al., 2020). It shows broad, but weak, molecular absorption features in comparison to what is seen in isolated planetary-mass objects (e.g., Manjavacas et al., 2020; Kirkpatrick et al., 2006). In Appendix H, we fit this spectrum using atmospheric models and show that no model performs better than a simple black-body highlighting the fact that models are not yet adapted for these young objects. One striking aspect of the spectrum is the large number of emission lines (also not present in models), whose properties are summarized in Table 2. These lines are commonly seen in young stellar objects and disks, and some of them have also been reported in PMCs (see details in Appendix E). Combined with the young age of the system ( Myr), these features strongly suggest that accretion is ongoing in V2376 Ori b. Interestingly, in the SED presented in Fig. 3, the filter F187N (Paschen ) shows an excess with respect to the reference broad filter F182M, suggesting that Pa emission, another accretion tracer, is also present. From the H and Pa emission, we estimate an accretion rate (see Appendix G). This value is comparable to those found for other PMCs (Eriksson et al., 2020; Zhou et al., 2014; Bowler et al., 2011; Santamaría-Miranda et al., 2017; Hashimoto et al., 2020) and PMOs (Mohanty et al., 2005; Fang et al., 2009; Joergens et al., 2013) with similar masses.
@@ 3.3 Extended structures
Fig. 2 C shows the sum of the integrated maps of the [O ii] doublet at 7320 and 7330 Å. These auroral lines indicate tenuous, ionized gas at a few thousand K. Considering that V2376 Ori is pc in front of the ONC, it seems unlikely that the gas is ionized by extreme UV ( eV) radiation from the ONC massive stars. Instead, it probably reflects ionization by local sources of extreme UV (e.g. from the accretion shocks). The extended [O ii] emission (Fig. 2) shows a structure that connects the companion to the star (“bridge”) and an arc-shaped structure to the West, reminiscent of a spiral arm. Spiral arms are now commonly seen in protoplanetary disks and can be related to the presence of planets within those disks (see e.g., Boccaletti et al., 2020; Haffert et al., 2019; Ginski et al., 2025; Hashimoto et al., 2020). However, we do not find any clear evidence of a large disk (70 au) around V2376 Ori in which the companion V2376 Ori b would be embedded. The absence of such a disk is also consistent with the fact that the large mass-ratio of the sources imply a rapid tidal truncation of a large circumstellar disk (Harris et al., 2012). Thus, it appears more likely that the observed “arc” and “bridge” are not related to the PMC being embedded in a large or circumbinary protoplanetary disk, but rather to some form of dynamical interaction between the two sources that results in mass transfer from their individual (small) accretion disks. Tidal interaction between young objects are shown to indeed create bridges and arcs similar to those observed in Fig. 2 C (Clarke & Pringle, 1993; Cabrit et al., 2006). A surprising aspect is that the structure is only seen in certain Oxygen lines, not in any of the other atomic lines reported in Table 2. This could reflect particular physical conditions leading to the excitation of these specific lines in the shocked gas or an over-abundance of oxygen with respect to other elements. Additional spectroscopy will be required to understand this.
@@ 4 Conclusion
Combined observations from JWST and the VLT have allowed us to confirm the presence of a planetary-mass companion with a mass of around the star V2376 Ori, located in the Orion D association. This result highlights the potential of JWST to identify new PMCs through NIRCam photometric imaging surveys of young associations. Our MUSE observations suggest the presence of a tidal interaction between the circumstellar and circumplanetary disks. Future observations, for instance with SPHERE at VLT to search for counterparts of the bridge and arc in dust scattering and JWST to better characterize the spectral types of the star and companion as well as disk properties are required. The upcoming ELT, with its more than tenfold improvement in angular and spectral resolution in the visible and near-IR, will be crucial to further characterize such dynamical interactions in young planetary systems.

### v2

# Observation of an accreting planetary-mass companion with signs of disc–disc interaction in Orion ††thanks: Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme ID 114.28JY

@@
Observation of an accreting planetary-mass companion with signs of disc–disc interaction in Orion ††thanks: Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme ID 114.28JY
Young ( Myr) planetary-mass companions (PMCs) provide valuable insights into the formation and early evolution of planetary systems. To date, only a dozen such objects have been identified through direct imaging. Using JWST/NIRCam observations towards the Orion Nebula, obtained as part of the PDRs4All Early Release Science program, we have identified a faint point source near the M-type star V2376 Ori, a member of Orion D, around 80 pc in the foreground of the Trapezium cluster of Orion and with an age of approximately Myr. Follow-up spectroscopic observations with the MUSE instrument on the VLT confirm that the source, V2376 Ori b, is indeed a young PMC. We fit the SED of V2376 Ori b and infer a mass of with evolutionary tracks. The MUSE spectrum reveals several accretion tracers. Based on the H line intensity, we estimate an accretion rate of 10, which is comparable to that of young PMCs such as PDS 70b. In addition, the MUSE data cube reveals extended emission in the [O ii] doublet at 7320 and 7330 Å, which is interpreted as evidence of a dynamical interaction between the two sources that potentially involves mass transfer between their individual accretion discs. These results demonstrate that JWST/NIRCam imaging surveys of young stellar associations can uncover new PMCs, which can then be confirmed and characterized through ground-based spectroscopic follow-up.
@@
brown dwarfs – exoplanets – stars: low-mass – accretion, accretion discs – direct imaging1 Introduction
Direct imaging has revealed key properties of planetary-mass companions (PMCs, , au; Currie et al. 2022). Only a dozen have been identified in young ( Myr) stellar clusters (e.g. Currie et al., 2013; Luhman et al., 2006; Schmidt et al., 2008; Bowler et al., 2014). Despite their rarity, young PMCs offer valuable constraints on planetary formation pathways (Kratter and Lodato, 2016; Schlaufman, 2018), especially in environments where external factors such as photoevaporation and dynamical interactions may play a role (Reiter and Parker, 2022; Planet formation environments collaboration et al., 2025). Expanding the sample of known young PMCs is critical but has been limited by the time-intensive nature of ground-based high-contrast imaging. JWST has recently demonstrated its potential to image both known and new PMCs as well as exoplanets (Boccaletti et al., 2024; Matthews et al., 2024; Carter et al., 2023; Mullally et al., 2024; Lagrange et al., 2025). In this letter, we report the discovery of a 7 Myr-old PMC in Orion, identified in JWST/NIRCam data from the PDRs4All Early Release Science program (PID 1288; Berné et al., 2022; Habart et al., 2024), and confirmed via follow-up spectroscopy with VLT/MUSE.
@@ 2 Source and data
V2376 Ori (, ) is a young star situated north-east of the Orion nebula cluster (ONC), catalogued as an M-type star by Hillenbrand (1997). In their catalogue of protoplanetary discs, Ricci et al. (2008) identified a faint source within of V2376 Ori, which the authors interpreted as a jet. This faint source was also reported by Habart et al. (2024) in the Paschen image obtained with the JWST. In these earlier studies, V2376 Ori was believed to be part of the ONC association. However, Gaia observations provide a parallax for V2376 Ori of mas, placing the source at a distance of pc, which is much closer than the ONC (390 pc, Maíz Apellániz et al., 2022). This implies that V2376 Ori is instead part of the Orion D association identified by Kounkel et al. (2018), a population of young stars (7-10 Myr) situated in front of the ONC. This foreground population has also been identified by Alves and Bouy (2012) as a region dominated by the NGC 1980 cluster, aged 4-5 Myr. Thus, we assume an age of Myr for V2376 Ori.
@@ 3 Results
Figure 1 presents the HST and JWST images of V2376 Ori obtained in 12 filters. The faint source previously reported by Ricci et al. (2008) and Habart et al. (2024) is clearly detected in all bands. Figure 2 shows the MUSE integral field spectroscopic observations of the system. It combines integrated maps and the spectrum of the companion over the full MUSE wavelength range. Panel 2A displays the white light image integrated over the 6908–9158 Å range. The companion source is clearly visible. Panel 2B shows the continuum-subtracted map (using the Background2D procedure described in Appendix C) of the forbidden [C i] 8727 line from neutral atomic carbon, where the companion is also detected. In addition, we find extended emission in the [O ii] doublet at 7320 and 7330 Å, which is shown in panel 2C as the sum of the integrated maps of both lines. A similar, although fainter, structure is also visible in the [O i] 6300 map. We analyse the relative astrometry of V2376 Ori and the nearby point source in Appendix B.1. We measure a angular separation of , corresponding to a projected separation of au when assuming a distance of pc. The relative position of the two objects has remained constant within instrumental uncertainties over 20 years, consistent with the source being gravitationally bound. We therefore identify it as a companion, hereafter referred to as V2376 Ori b. Panel 2D presents the MUSE spectrum of this companion.
3.1 Temperature, luminosity, and mass of V2376 Ori b
Figure 3 presents the HST and JWST spectral energy distribution (SED) of V2376 Ori b, obtained as explained in Appendix C, within the dashed red circular aperture of 0.08” radius, shown in Fig. 1 (G). We fitted a black-body using a grid of temperatures ranging from 1000 K to 3000 K with a step of 100 K, and a scaling coefficient, with values ranging from to . The best-fit parameters are those that minimize the root mean square error (RMSE). The error bars were determined considering the parameters range for which the RMSE is within 1.5 times the minimum RMSE. This yielded K. The derived black-body luminosity is , where the uncertainty arises from the errors on both the temperature and the distance to the source. As the system is situated in a young association, extinction from interstellar or circumstellar dust may affect the SED fitting. To investigate this, we used the extinction curve of Weingartner and Draine (2001) with parametrized by the visual extinction, , in the fit. The best fit gives =0 suggesting a negligible effect of extinction. In Fig. 3 we also test the use of the BT-Cond model (Allard et al., 2001) to fit the SED; however, this gives poorer fits to the SED (see also the detailed comparison in Appendix H). Figure 6 shows the position of V2376 Ori b on the evolutionary tracks from Baraffe et al. (2003). Considering the obtained luminosity, effective temperature, and an age of 73 Myr, this yields a mass in the range of 10-30 . The derived physical properties of V2376 Ori b are summarized in Table 1. Figure 7 shows that these properties are similar to those found for other directly imaged PMCs and giant exoplanets.
@@ 3.2 MUSE spectroscopy of the V2376 Ori system
Figure 10 presents the MUSE spectrum of V2376 Ori (see details on the extraction method in Appendix E). This spectrum is found to be compatible with that of an M7.5-9 spectral type (Appendix F). Figure 2 (D) presents the MUSE spectrum for V2376 Ori b. It shows smooth continuum emission, with an increase towards the red and numerous emission lines. This spectrum is similar to that of the MUSE spectrum of the 12 MJup PMC Delorme 1 (AB; Eriksson et al. 2020). It shows broad, but weak, molecular absorption features in comparison to what is seen in isolated planetary-mass objects (e.g. Manjavacas et al., 2020; Kirkpatrick et al., 2006). In Appendix H we fit this spectrum using atmospheric models and show that no model performs better than a simple black-body, highlighting the fact that models are not yet adapted for these young objects. One striking aspect of the spectrum is the large number of emission lines (also not present in models), whose properties are summarized in Table 2. These lines are commonly seen in young stellar objects and discs, and some of them have also been reported in PMCs (see details in Appendix E). Combined with the young age of the system ( Myr), these features strongly suggest that accretion is ongoing in V2376 Ori b. Interestingly, in the SED presented in Fig. 3, the filter F187N (Paschen ) shows an excess with respect to the reference broad filter F182M, suggesting that Pa emission, another accretion tracer, is also present. From the H and Pa emission, we estimate an accretion rate (see Appendix G). This value is comparable to those found for other PMCs (Eriksson et al., 2020; Zhou et al., 2014; Bowler et al., 2011; Santamaría-Miranda et al., 2017; Hashimoto et al., 2020) and PMOs (Mohanty et al., 2005; Fang et al., 2009; Joergens et al., 2013) with similar masses.
@@ 3.3 Extended structures
Figure 2 C shows the sum of the integrated maps of the [O ii] doublet at 7320 and 7330 Å. These auroral lines indicate tenuous, ionized gas at a few thousand Kelvin. Considering that V2376 Ori is pc in front of the ONC, it seems unlikely that the gas is ionized by extreme UV ( eV) radiation from the ONC massive stars. Instead, it probably reflects ionization by local sources of extreme UV (e.g. from the accretion shocks). The extended [O ii] emission (Fig. 2) shows a structure that connects the companion to the star (‘bridge’) and an arc-shaped structure to the west, reminiscent of a spiral arm. Spiral arms are now commonly seen in protoplanetary discs and can be related to the presence of planets within those discs (see e.g. Boccaletti et al., 2020; Haffert et al., 2019; Ginski et al., 2025; Hashimoto et al., 2020). However, we do not find any clear evidence of a large disc (70 au) around V2376 Ori in which the companion V2376 Ori b would be embedded. The absence of such a disc is also consistent with the fact that the large mass-ratio of the sources imply a rapid tidal truncation of a large circumstellar disc (Harris et al., 2012). Thus, it appears more likely that the observed ‘arc’ and bridge are not related to the PMC being embedded in a large or circumbinary protoplanetary disc, but rather to some form of dynamical interaction between the two sources that results in mass transfer from their individual (small) accretion discs. Tidal interaction between young objects are shown to indeed create bridges and arcs similar to those observed in Fig. 2 C (Clarke and Pringle, 1993; Cabrit et al., 2006). A surprising aspect is that the structure is only seen in certain oxygen lines, not in any of the other atomic lines reported in Table 2. This could reflect particular physical conditions leading to the excitation of these specific lines in the shocked gas or an over-abundance of oxygen with respect to other elements. Additional spectroscopy will be required to understand this.
@@ 4 Conclusion
Combined observations from JWST and the VLT have allowed us to confirm the presence of a PMC with a mass of around the star V2376 Ori, located in the Orion D association. This result highlights the potential of JWST to identify new PMCs through NIRCam photometric imaging surveys of young associations. Our MUSE observations suggest the presence of a tidal interaction between the circumstellar and circumplanetary discs. Future observations, for instance with SPHERE at VLT, to search for counterparts of the bridge and arc in dust scattering and JWST to better characterize the spectral types of the star and companion as well as disc properties are required. The upcoming ELT, with its more than tenfold improvement in angular and spectral resolution in the visible and near-IR, will be crucial to further characterize such dynamical interactions in young planetary systems.